OpenStudy (anonymous):

Evaluate the limit using the identity a3−b3=(a−b)(a2+ab+b2) ... lim x goes to 7 ... (x^3-343) / (x-7)

OpenStudy (anonymous):

\[\frac{x^3-343}{x-7}=\frac{(x-7)(x^2+7x+49)}{x-7}=\cdots\]

OpenStudy (anonymous):

You're a genius